x+y+z+xy+yz+zx = 89 find the no of integral solutions? 

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
this ques is wrong. correct ques should be x+y+z+xy+yz+zx+xyz = 89
we can then solve it as follows:
add 1 on both sides
1+x+y+z+xy+yz+zx+xyz = 90
or (1+x)(1+y)(1+z)= 90
now this turns into a simple problem of finding the number of ways to factorise 90 into 3 factors.
kindly approve :)